# Tasa: Thermal-aware 3D-Stacked Architecture Design with Bandwidth Sharing for LLM Inference

### v1

@@ I Introduction
Large language models (LLMs), such as GPTs [1] and Llama [2], have exhibited extraordinary ability for diverse natural language processing applications, including conversational agents, text generation, summarization, etc. LLM inference consists of two operation stages, i.e. prefill and decoding. In the prefill stage, the LLM takes the prompt tokens as input to generate the first output token. After that, the LLM begins autoregressive decoding, using the last generated tokens and previous context as input to predict the next token in each iteration. Prefill and decoding of LLMs pose stringent high requirements for computation performance and memory bandwidth, respectively.
@@
Prior studies [3, 4, 5, 6] show that the sequential autoregressive decoding is the latency bottleneck of LLM inference due to its memory-intensive characteristics and the limited memory bandwidth. For example, NVIDIA A100 GPUs [7] have high computing performance (312 TFLOPS) but limited off-chip DRAM bandwidth (1555 GB/s). To address the memory bandwidth bottleneck, processing-in-memory (PIM) technique is a promising approach, which integrates processing units near DRAM banks and utilizes the high internal bank-level memory bandwidth. In recent years, DRAM vendors presented PIM designs with enhanced general matrix-vector multiplication (GEMV) performance, such as GDDR-PIM from SK Hynix [8], HBM-PIM and LPDDR-PIM from Samsung [9, 10, 11]. Furthermore, prior works [3, 5, 4] have explored GPU-PIM heterogeneous systems to optimize LLM inference, in which memory-intensive attention layers are offloaded to PIM while the compute-intensive fully connected layers are executed on GPU. However, these architectures are only efficient for scenarios with high operation intensity (Op/B) and does not perform well across a wide range of Op/B.
@@
The limitations of existing solutions motivate architecture explorations with higher memory bandwidth, such as 3D-stacked logic-to-DRAM architecture [12, 13, 14, 15]. In NVIDIA’s vision for future accelerator design, the 3D-stacked architecture will be employed to keep scale compute [16]. Such 3D-stacked architecture vertically stacks DRAM dies on top of a logic die via hybrid bonding (HB), where the DRAM and log dies utilize independent process technology. The state-of-the-art stacking technology has also developed high-density hybrid bonding vias and fine pitch mini-TSVs [17, 18] for higher IO parallelism and memory bandwidth.
@@
Although 3D-stacked architecture has promising attributes, it is faced with serious thermal issues for applications such as LLM inferences. When there are more dies in vertical direction, the power density will increase accordingly and the thermal issue will have a significant impact on reliability, efficiency, and longevity of 3D stacked chips. As reported in [19], thermal failures account for 55% of all issues encountered in 3D stacked chips. Based on our analysis, the thermal issue becomes more severe for high-performance designs with larger 3D stacked areas, significantly impacting the computing performance, i.e. similar to dark silicon [20]. Due to these thermal issues, the 3D-stacked designs have restricted power constraints and require more careful architectural-level thermal considerations compared to 2D architectures.
@@ II-A LLMs and 3D Architectures
Modern LLM inference consists of two distinct stages: the prefill and decoding. During prefill, the model takes the prompt sequence as input and generates the first token and KV cache for decoding. In the decoding stage, the model iteratively takes the last output token and previous KV cache as input to generate a new token and update the KV cache. LLMs are structured based on Transformer blocks [21], which consists of a multi-head attention (MHA) module and a feed-forward network (FFN), accompanied by layer normalization and residual operators. Due to the sequential nature of autoregressive decoding, the decoding accounts for most of the end-to-end inference latency. For example, the decoding stage consumes more than 96% of the total execution time in the GPT-3 175B model for 32 input tokens and output tokens [3].
To enhance the LLM inference throughput, emerging 3D-stacked architecture with HB is a promising solution. The architecture stacks multi-DRAM dies on top of a logic die with HB and mini-TSV technology. The dummy HB is also employed across layers to reduce thermal stress and temperature gap [14]. Fig. 1 shows a typical 3D design topped with thermal interface material (TIM) and a heat sink. For 3D-stacked DRAM, the array blocks are distributed with fine granularity. Each block is an independent memory channel and is connected to its local logic core by HB. The high density HB via (110,00/ with 3um pitch) and 1.5um fine pitch mini-TSV unlocks extremely high I/O parallelism. The direct bonding process significantly reduces parasitic resistance and capacitance, resulting in much lower logic-to-memory interface power consumption than HBM. For example, SeDRAM [17] is 3D connected to logic die through HB and achieves 34GB ps/Gbit bandwidth and 0.88 pJ/bit efficiency, which is 6.2 higher bandwidth and 4.35 better efficiency than HBM3 [22]. Prior works have exploited 3D-stacked architectures in various scenarios. Niu et al. [13] presented a 3D-stacked processing-near-memory engine for the memory-bound recommendation system. Yue et al. [15] accelerate vision AI models with 3D-stacked DRAM.
@@ II-B Thermal Considerations for 3D ICs
Thermal is always a challenging issue for high-performance computing, particularly in 3D chips. The vertically stacked dies not only enhance power density but also introduce higher thermal resistance. The larger distance between logic die and the heat sink further diminishes the efficiency of the cooling effect, resulting in poor heat dissipation. These thermal limitations restrict the achievable peak performance and memory bandwidth, ultimately degrading the advantages of the 3D-stacked architecture. To improve thermal dissipation, researchers have investigated various structures and materials, alongside innovative cooling methods to address the heat dissipation in 3D chips [23, 24].
In addition to the passive cooling solutions, diverse research has explored dynamic thermal managements (DTMs) based on runtime workloads [25, 26, 27], such as dynamic voltage-frequency scaling (DVFS) and thread migration in CPUs. These thermal management schemes can dynamically regulate chip operation conditions and actively mitigate overheating and local hot spots. Furthermore, a few studies also optimize the memory access and data storage layout in DRAM based on the memory access pattern [28, 29, 30] to reduce DTM overhead.
@@ III Thermal Analysis of 3D Architecture
To well understand the thermal challenges for 3D-stacked architectures, we perform comprehensive profiling and analysis of temperature distribution in 3D architecture. As shown in Fig. 2, we experiment with a 4-layer stacked memory with 8 TB/s bandwidth and a logic die which can be abstracted into memory, communication unit, and the computation array. For the computation array, a GPU-like architecture [7] is adopted with 48 cores to achieve a performance of 196 TFLOPS, in which each core features two 3232 systolic arrays and occupies an area of 4mm2.5mm. We evaluate the thermal profile for GPT-3 66B inference with a batch size of 32 using HotSpot 6.0 simulator [31]. The detailed hardware simulation configurations are provided in Table II.
@@ V-A Dataflow and Mapping
Fig. 8 (a) shows the platform architecture using Tasa for LLM inference. To accommodate to the LLM model size, multiple 3D-stacked Tasa architecture can be integrated as a inference platform. The interconnect between Tasa devices and host CPU adopts standard interconnect such as PCIe. Fig. 8 (b) top shows our dataflow across Tasa. Simiar as the dataflow in heterogeneous systems [3], in our case, P-cores process computation-intensive FC layers in QKV generation and projection while the memory-intensive attention layers are offloaded to E-cores. It adopts the head-level pipeline, in which the FC layers for QKV generation and projection are pipelined with attention layers to reduce overall latency. To ensure high parallelism, tensor parallelism is implemented across the devices to enhance the performance. Within each device, the model parameters are evenly partitioned across cores at the tensor level to avoid remote memory access. The weight for QKV generation and projection is distributed within P-cores, while the weight for FFN is distributed in both P-cores and E-cores. The KV cache is all stored in E-cores.
@@ VI-A Experimental Setup
Workload: We set LLaMA-65B [32], GPT-3 66B [1] as the target model with FP16/BF16 data type. We conduct our evaluation using real-world conversational requests sampled from ShareGPT, which is a set of conversations taken from the real-world user log of ChatGPT. We evaluate the performance of stable batch sizes by dynamically adding the next request whenever an inference completes. For varying queries per second (QPS) situations, we generate request arrivals through a Poisson process to simulate realistic traffic patterns.
@@
Simulation: We developed a simulator framework by a combination of AttAcc [3], Timeloop [33], Accelergy [34], Ramulator 2.0 [35], and HotSpot 6.0 [31] for evaluations. To evaluate the performance, we assess the performance of each operator using Timeloop, with DRAM accesses offloaded to Ramulator. To evaluate the area and energy, we extend Accelergy for 7nm process (ASAP7) [36] and the 3D-stacked DRAM evaluation. we modified FinCACTI with the published information of SRAM to model 7nm SRAM. The energy consumption of stacked DRAM refers to the latest data from [18], while the area of the memory controller is estimated from [13]. Finally, we augment the AttAcc framework with the above results for end-to-end inference evaluation and power trace generation. For thermal evaluation, we utilize HotSpot 6.0 to analyze the temperature with the floorplan and power traces, with detailed configuration in Table. II.
@@ VII Discussions
Furthermore, other optimization techniques such as thermal-aware floorplanning are orthogonal to our approach. For example, prior research has investigated floorplanning method for many-core 3D systems [37] to refine both horizontal and vertical layouts, which can also be applied to augment Tasa. The fundamental contributions of Tasa lie in introducing heterogeneous architecture optimized for LLM inference and demonstrating its superiority over homogeneous designs in both thermal and inference metrics through comprehensive evaluations, which enable further research on floorplanning optimization for heterogeneous 3D-stacked architectures.

### v3

@@ I Introduction
Large language models (LLMs), such as GPTs [gpt] and Llama [llama], have exhibited extraordinary ability for diverse natural language processing applications, including conversational agents, text generation, summarization, etc. LLM inference consists of two operation stages, i.e. prefill and decoding. In the prefill stage, the LLM takes the prompt tokens as input to generate the first output token. After that, the LLM begins autoregressive decoding, using the last generated tokens and previous context as input to predict the next token in each iteration. Prefill and decoding of LLMs pose stringent high requirements for computation performance and memory bandwidth, respectively.
@@
Prior studies [attacc_2024, specpim_2024, neupims_2024, duplex] show that the sequential autoregressive decoding is the latency bottleneck of LLM inference due to its memory-intensive characteristics and the limited memory bandwidth. For example, NVIDIA A100 GPUs [a100] have high computing performance (312 TFLOPS) but limited off-chip DRAM bandwidth (1555 GB/s). To address the memory bandwidth bottleneck, processing-in-memory (PIM) technique is a promising approach, which integrates processing units near DRAM banks and utilizes the high internal bank-level memory bandwidth. In recent years, DRAM vendors presented PIM designs with enhanced general matrix-vector multiplication (GEMV) performance, such as GDDR-PIM from SK Hynix [gddr_pim], HBM-PIM and LPDDR-PIM from Samsung [hbm_pim, lpddr_pim, samsung_PIM]. Furthermore, prior works [attacc_2024, neupims_2024, specpim_2024] have explored GPU-PIM heterogeneous systems to optimize LLM inference, in which memory-intensive attention layers are offloaded to PIM while the compute-intensive fully connected layers are executed on GPU. However, these architectures are only efficient for scenarios with high operation intensity (Op/B) and does not perform well across a wide range of Op/B.
@@
The limitations of existing solutions motivate architecture explorations with higher memory bandwidth, such as 3D-stacked logic-to-DRAM architecture [tetris_3d, alibaba_3d, asscc_3d, iscca24_3d]. In NVIDIA’s vision for future accelerator design, the 3D-stacked architecture will be employed to keep scale compute [nvidia_3d]. Such 3D-stacked architecture vertically stacks DRAM dies on top of a logic die via hybrid bonding (HB), where the DRAM and log dies utilize independent process technology. The state-of-the-art stacking technology has also developed high-density hybrid bonding vias and fine pitch mini-TSVs [sedram, sedram2] for higher IO parallelism and memory bandwidth.
@@
Although 3D-stacked architecture has promising attributes, it is faced with serious thermal issues for applications such as LLM inferences. When there are more dies in vertical direction, the power density will increase accordingly and the thermal issue will have a significant impact on reliability, efficiency, and longevity of 3D stacked chips. As reported in [thermal_failure], thermal failures account for 55% of all issues encountered in 3D stacked chips. Based on our analysis, the thermal issue becomes more severe for high-performance designs with larger 3D stacked areas, significantly impacting the computing performance, i.e. similar to dark silicon [dark_silicon]. Due to these thermal issues, the 3D-stacked designs have restricted power constraints and require more careful architectural-level thermal considerations compared to 2D architectures.
@@ II-A LLMs and 3D Architectures
Modern LLM inference consists of two distinct stages: the prefill and decoding. During prefill, the model takes the prompt sequence as input and generates the first token and KV cache for decoding. In the decoding stage, the model iteratively takes the last output token and previous KV cache as input to generate a new token and update the KV cache. LLMs are structured based on Transformer blocks [attention], which consists of a multi-head attention (MHA) module and a feed-forward network (FFN), accompanied by layer normalization and residual operators. Due to the sequential nature of autoregressive decoding, the decoding accounts for most of the end-to-end inference latency. For example, the decoding stage consumes more than 96% of the total execution time in the GPT-3 175B model for 32 input tokens and output tokens [attacc_2024].
To enhance the LLM inference throughput, emerging 3D-stacked architecture with HB is a promising solution. The architecture stacks multi-DRAM dies on top of a logic die with HB and mini-TSV technology. The dummy HB is also employed across layers to reduce thermal stress and temperature gap [asscc_3d]. Fig. 1 shows a typical 3D design topped with thermal interface material (TIM) and a heat sink. For 3D-stacked DRAM, the array blocks are distributed with fine granularity. Each block is an independent memory channel and is connected to its local logic core by HB. The high density HB via (110,00/ with 3um pitch) and 1.5um fine pitch mini-TSV unlocks extremely high I/O parallelism. The direct bonding process significantly reduces parasitic resistance and capacitance, resulting in much lower logic-to-memory interface power consumption than HBM. For example, SeDRAM [sedram] is 3D connected to logic die through HB and achieves 34GB ps/Gbit bandwidth and 0.88 pJ/bit efficiency, which is 6.2 higher bandwidth and 4.35 better efficiency than HBM3 [hbm3]. Prior works have exploited 3D-stacked architectures in various scenarios. Niu et al. [alibaba_3d] presented a 3D-stacked processing-near-memory engine for the memory-bound recommendation system. Yue et al. [iscca24_3d] accelerate vision AI models with 3D-stacked DRAM.
@@ II-B Thermal Considerations for 3D ICs
Thermal is always a challenging issue for high-performance computing, particularly in 3D chips. The vertically stacked dies not only enhance power density but also introduce higher thermal resistance. The larger distance between logic die and the heat sink further diminishes the efficiency of the cooling effect, resulting in poor heat dissipation. These thermal limitations restrict the achievable peak performance and memory bandwidth, ultimately degrading the advantages of the 3D-stacked architecture. To improve thermal dissipation, researchers have investigated various structures and materials, alongside innovative cooling methods to address the heat dissipation in 3D chips [thermal_review_1, thermal_review_2].
In addition to the passive cooling solutions, diverse research has explored dynamic thermal managements (DTMs) based on runtime workloads [thermal_cpu, thermal_dvfs, thermal_dvfs2], such as dynamic voltage-frequency scaling (DVFS) and thread migration in CPUs. These thermal management schemes can dynamically regulate chip operation conditions and actively mitigate overheating and local hot spots. Furthermore, a few studies also optimize the memory access and data storage layout in DRAM based on the memory access pattern [thermal_dram0, thermal_dram1, thermal_dram2] to reduce DTM overhead.
@@ III Thermal Analysis of 3D Architecture
To well understand the thermal challenges for 3D-stacked architectures, we perform comprehensive profiling and analysis of temperature distribution in 3D architecture. As shown in Fig. 2, we experiment with a 4-layer stacked memory with 8 TB/s bandwidth and a logic die which can be abstracted into memory, communication unit, and the computation array. For the computation array, a GPU-like architecture [a100] is adopted with 48 cores to achieve a performance of 196 TFLOPS, in which each core features two 3232 systolic arrays and occupies an area of 4mm2.5mm. We evaluate the thermal profile for GPT-3 66B inference with a batch size of 32 using HotSpot 6.0 simulator [HotSpot6_2015]. The detailed hardware simulation configurations are provided in Table II.
@@ V-A Dataflow and Mapping
Fig. 8 (a) shows the platform architecture using Tasa for LLM inference. To accommodate to the LLM model size, multiple 3D-stacked Tasa architecture can be integrated as a inference platform. The interconnect between Tasa devices and host CPU adopts standard interconnect such as PCIe. Fig. 8 (b) top shows our dataflow across Tasa. Simiar as the dataflow in heterogeneous systems [attacc_2024], in our case, P-cores process computation-intensive FC layers in QKV generation and projection while the memory-intensive attention layers are offloaded to E-cores. It adopts the head-level pipeline, in which the FC layers for QKV generation and projection are pipelined with attention layers to reduce overall latency. To ensure high parallelism, tensor parallelism is implemented across the devices to enhance the performance. Within each device, the model parameters are evenly partitioned across cores at the tensor level to avoid remote memory access. The weight for QKV generation and projection is distributed within P-cores, while the weight for FFN is distributed in both P-cores and E-cores. The KV cache is all stored in E-cores.
@@ VI-A Experimental Setup
Workload: We set LLaMA-65B [llama-65B], GPT-3 66B [gpt] as the target model with FP16/BF16 data type. We conduct our evaluation using real-world conversational requests sampled from ShareGPT, which is a set of conversations taken from the real-world user log of ChatGPT. We evaluate the performance of stable batch sizes by dynamically adding the next request whenever an inference completes. For varying queries per second (QPS) situations, we generate request arrivals through a Poisson process to simulate realistic traffic patterns.
@@
Simulation: We developed a simulator framework by a combination of AttAcc [attacc_2024], Timeloop [timeloop], Accelergy [accelergy], Ramulator 2.0 [ramulator2_2023], and HotSpot 6.0 [HotSpot6_2015] for evaluations. To evaluate the performance, we assess the performance of each operator using Timeloop, with DRAM accesses offloaded to Ramulator. To evaluate the area and energy, we extend Accelergy for 7nm process (ASAP7) [asap7_2016] and the 3D-stacked DRAM evaluation. we modified FinCACTI with the published information of SRAM to model 7nm SRAM. The energy consumption of stacked DRAM refers to the latest data from [sedram2], while the area of the memory controller is estimated from [alibaba_3d]. Finally, we augment the AttAcc framework with the above results for end-to-end inference evaluation and power trace generation. For thermal evaluation, we utilize HotSpot 6.0 to analyze the temperature with the floorplan and power traces, with detailed configuration in Table. II.
@@ VII Discussions
Furthermore, other optimization techniques such as thermal-aware floorplanning are orthogonal to our approach. For example, prior research has investigated floorplanning method for many-core 3D systems [floorplan1] to refine both horizontal and vertical layouts, which can also be applied to augment Tasa. The fundamental contributions of Tasa lie in introducing heterogeneous architecture optimized for LLM inference and demonstrating its superiority over homogeneous designs in both thermal and inference metrics through comprehensive evaluations, which enable further research on floorplanning optimization for heterogeneous 3D-stacked architectures.